OpenStudy (anonymous):

The ratio for the lengths of radii of similar cylinders is 6 : 4. What is the ratio of their volumes? I got 27:4

OpenStudy (anonymous):

Am I correct

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

hm no. what's the equation for volume of cylinders?

OpenStudy (anonymous):

pi(rsquared)(h)

OpenStudy (anonymous):

good, now you plug 6 into that equation. pi 6 squared h = 36 pi h

OpenStudy (anonymous):

then, find volume of the other cylinder. Plug 4 into that equation pi 4 squared (h) = 16 pi h

OpenStudy (anonymous):

Now, you know that ratio can be written as fractions,right? so 36pi h : 16 pi h \[= \frac{ 36 \pi h }{ 16 \pi h }\]

OpenStudy (anonymous):

now you can cross out pi and h on top and bottom

OpenStudy (anonymous):

\[\large = \frac{ 36 }{ 16 } = \frac{ 9 }{ 4 }\] so the ratio is 9 : 4

OpenStudy (anonymous):

oh I get it thanks

OpenStudy (anonymous):

anytime :D

OpenStudy (anonymous):

can you help me with this one I don't get it. If the scale factor between two similar solids is 2:5, then the ratio of their lateral areas is 4:10

OpenStudy (anonymous):

is this a true/false question?

OpenStudy (anonymous):

yeah it it.

OpenStudy (anonymous):

I thought it was false... just a guess though

OpenStudy (anonymous):

I found a good link :) hold on

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

Took the pic, same with lateral area/surface area.

OpenStudy (anonymous):

so, if you squared 2:5 the answer you get is 4 : 25

OpenStudy (anonymous):

oh so lateral area is the same as surface area.

OpenStudy (anonymous):

no, you use this equation for both surface area and lateral area. They are not the same

OpenStudy (anonymous):

oh okay thanks

OpenStudy (anonymous):

Lateral areas does not include base's area. Surface area includes BOTH sides + base areas.

OpenStudy (anonymous):

You're welcomed :)